OpenStudy (ahvei):

Mr. Cooper wants to know how many quarts of paint he will need to repaint the office. The dimensions of the office are 16 ft. by 12 ft. by 10 ft. The floor, ceiling and door do not need to be painted. The door is 3 ft. by 8 ft. One quarter of paint covers 100 square feet. How many quarts will they need if only full quarts can be purchased? Show your work!......I really need help on this one by 4:45 please answer fast if you know message me if you need me to answer one of your questions I will return the favor..... thanks in advance muah!!!!

OpenStudy (anonymous):

To find out how much paint is needed, we need to figure out the total area of the entire office then subtract from that the area of the sections not being painted, which include the ceiling, floor, and the door. The office is a rectangular shape, and the formula for area of a rectangle is Area = 2(wh + wl + lh), so let’s plug in the values from the problem: Area = 2 (10)(12) + (10)(16) + (16)(12). Now we need to simplify: Area = 2 (120 + 160 + 192) … Area = 2 (472) … Area = 944 ft.2 Now let’s find out the areas of the sections that are not being painted. The floor measures 10 ft x 16 ft, which is 320 ft. The ceiling is the same. The door measures 8 ft x 3 ft, so the area is 24 ft.2 The total area that is not being painted is 344 ft.2, so let’s subtract this from our total area of 944 ft.2: 944 – 344 = 600 ft.2 The problem says that one quart of paint covers 100 square feet, so we need to divide 600 by 100 to get 6. Mr. Cooper will need 6 full quarts of paint to cover his office.

OpenStudy (ahvei):

@I_Always_Smiling , @srgntb0$$